OpenStudy (anonymous):

Simplify. Remember to use absolute-value notation when necessary. If a root cannot be simplified, state this. it have a check sign√(a+3)^2

OpenStudy (anonymous):

\[\sqrt{(a+3)^2}=a+3\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

wht

OpenStudy (anonymous):

\[\sqrt{(x+3)^2}=|a+3|\]

OpenStudy (anonymous):

√(a+3)^2 =(a+3)

OpenStudy (anonymous):

\[\sqrt{(a+3)^2} = \left| a+3 \right|\]

OpenStudy (anonymous):

this . Remember to use absolute-value notation when necessary was your hint